OpenStudy (anonymous):

how to find the domain of a function square root 1-x

OpenStudy (mathmate):

do you mean \[f(x)=\sqrt{1-x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The domain of a function is the values of the independent variable (in this case x), of which the function is defined. In other words, the values that aren't undefined. Can you think of some values for x in the function\[f(x)=\sqrt{x-1}\] where it is not defined? (Hint: the square root of a negative number is not defined)

OpenStudy (anonymous):

would it be all real numbers except 1

OpenStudy (anonymous):

Not quite. What is zero minus one. How about negative ten minus one. And negative one thousand minus one?

OpenStudy (anonymous):

Oh I'm sorry, I have the function wrong..... Think about one minus two. Or one minus ten. And one minus one thousand.

OpenStudy (anonymous):

so what would square root 1-x be

OpenStudy (anonymous):

It would be easier to understand if you would start using the equation function of the forum. \[f(x)=\sqrt{1-x}\]This is a mathematical function. When you define the domain for this function, you are going to be using an interval. That means a range of numbers. For example, the domain of the function\[g(x)=4/(5-x)\]is all real numbers except for the value 5, because division by zero is undefined.